Find the following integral, given that S Sadar f(x)dx = 9 and ce)?dx = 13 a lisaslan - /sear -
Answer #1

solution

given that   \int_{a}^{b}f(x)dx=9 and \int_{a}^{b}(f(x))^2dx=13

\therefore \int_{a}^{b}(f(x))^2dx-[\int_{a}^{b}f(x)dx]^2=13-(9)^2

  =13-81

  =-68 answer